A die contains the numbers one through six. If the die is rolled twice, what is the probability of it landing on five in two consecutive tosses?
(A) 2/6 (B) 5/6 (C) 1/36 (D) 10/36 (E) 25/36

Answer by stanbon(75887)

A die contains the numbers one through six. If the die is rolled twice, what is the probability of it landing on five in two consecutive tosses?
---
The toss results are independent.
P(5 and 5) = (1/6)(1/6) = 1/36
---------
Cheers,
Stan H.
==========
(A) 2/6 (B) 5/6 (C) 1/36 (D) 10/36 (E) 25/36